OpenStudy (anonymous):

Convert 38degrees F to degrees celsius. If necessary, round your answer to the nearest tenth of a degree. Here are the formulas. C=5/9(F-32) F=9/5 C+32

OpenStudy (amistre64):

what is 38-32?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

plug in the f which is 38 . intot he first equation:

OpenStudy (amistre64):

good, and 6*5?

OpenStudy (amistre64):

liliy has a typo

OpenStudy (anonymous):

ooopppss! thanks

OpenStudy (anonymous):

5/9(38-32)

OpenStudy (anonymous):

30

OpenStudy (amistre64):

good, and the last part is 30/9 = C; whats 30/9?

OpenStudy (anonymous):

3.3

OpenStudy (amistre64):

ill trust you on that one

OpenStudy (anonymous):

lol so the answer is ? 3.3?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Omg! that was easy, Thank yu

OpenStudy (amistre64):

yw :)

OpenStudy (anonymous):

Whats 17.7 rounded to the nearest tenth of a degree?

OpenStudy (anonymous):

.7 is the tenth

OpenStudy (anonymous):

so the answer is 18?

OpenStudy (anonymous):

no , u keep it the decimal . a TENTH is 1/10 of a number. so its the first place after the decimal...

OpenStudy (anonymous):

I don't get it...Please explain?

OpenStudy (anonymous):

the number 17.7 : the 1 is the tens place, 7 is hte ones place, .7 is the TENTHS place. so when it asks for the tenth place, it wants the number up until that place: which owuld be 17.7

OpenStudy (anonymous):

Okay Thank you.. But I just put that in and it says its wrong

OpenStudy (anonymous):

hmmmm.... does it want hte TENs or TENTHs place?

OpenStudy (anonymous):

the tenths place

OpenStudy (anonymous):

that is strange. where are you putting this into?

OpenStudy (anonymous):

a math program for my math class

OpenStudy (anonymous):

well the number after the decimal is the tenths place.

OpenStudy (anonymous):

yes, aren't you suppose to round up because the number is 7?

OpenStudy (anonymous):

but you round to that number.. wats the number after the 7?

OpenStudy (anonymous):

its alot of 7's after

OpenStudy (anonymous):

oh, well then yes, you round up. so your answer should be 17.8

OpenStudy (anonymous):

Nope its wrong...but thank yu anyway...

OpenStudy (anonymous):

oh. well thats awkward. lol sorry

OpenStudy (anonymous):

lOl it said the answer was 147.2

OpenStudy (anonymous):

thats not even possible. F is always a bigger number than C. if F is 38, C is gonna be really small....

OpenStudy (anonymous):

Thats wat Im saying, idk... but this is how it has the formula f=9/5c+32 =9/5(64)+32 =147.2

OpenStudy (anonymous):

where did it get 64 from????

OpenStudy (anonymous):

i had to convert 64 degrees C to degrees F

OpenStudy (anonymous):

oy vey.

OpenStudy (anonymous):

Idk ma...

OpenStudy (anonymous):

But (64)+32 = 96 so I dn't get it

OpenStudy (anonymous):

but u need to multiply by 9/5

OpenStudy (anonymous):

but that gives me 172.8 or did i do it wrong?

OpenStudy (anonymous):

u did it wrong.... it shoudl be 147.2

OpenStudy (anonymous):

you are right...my bad! I got it now i have to use a scientific calculator

OpenStudy (anonymous):

thank yu btw....

OpenStudy (anonymous):

np.